OpenStudy (anonymous):

find exact value of expression without a calculator. 1-tan^2 65 +csc^2 25

OpenStudy (amoodarya):

csc^2 25=1+cot^2 25

OpenStudy (amoodarya):

25+65=90 so sin(25)=cos(65) or tan(25)=cot(65) ...

OpenStudy (anonymous):

Replace and you find that your expressions boils down to 2